# Pre-trained Prompt-driven Community Search

### v1

@@ IV-A Settings
(1) An Unsupervised community search algorithm includes the M method.
@@
TABLE IV: Results of M, SLSS, ComAF, SLRL, and PPCS. The best results are highlighted in bold, and the second-best results are underlined.
Datasets Metrics M SLSS CommunityAF SLRL PPCS
@@
(2) Semi-supervised community search algorithms include SLSS, CommunityAF, and SLRL.
  • SLSS [24]: This method is the first semi-supervised community search algorithm, which proposes a graph kernel-based community structure similarity measurement.
  • CommunityAF [28]: This method is a semi-supervised community search model consisting of three components: a GNN-based feature extraction module, an autoregressive flow-based community generation module, and a scoring module.
@@ IV-B Results
IV-B1 Comparison with semi-supervised community search algorithms
We compared the results of M, SLSS, CommunityAF, SLRL, and PPCS. Table IV presents the experimental results for five datasets.
Experimental results indicate that PPCS outperforms M, SLSS, CommunityAF, and SLRL in terms of overall performance. On the Amazon, DBLP, Twitter, and YouTube datasets, PPCS achieves higher F-score and Jaccard values compared to the other methods, except the LiveJournal dataset, where its performance is slightly inferior. In general, semi-supervised community search algorithms (SLSS, CommunityAF, SLRL, and PPCS) tend to surpass the unsupervised method M, as they leverage known community information as prior knowledge, thereby guiding the model to identify community structures more accurately. However, the M method exhibits slightly better performance on the LiveJournal dataset, because the community structure in this dataset aligns more closely with the modeling assumptions underlying the M method.
On most datasets other than LiveJournal, PPCS outperforms SLRL. This is primarily due to PPCS’s ability to learn structural characteristics of local structures through the node encoding component, which enhances the model’s generalization capacity. Additionally, PPCS incorporates a prompt mechanism to generate query-specific community predictions, thereby improving prediction accuracy. In contrast, SLRL performs relatively poorly on the LiveJournal dataset due to its inability to learn effective features with limited labeled data, whereas PPCS maintains strong predictive performance even with few samples. PPCS outperforms CommunityAF because it utilizes query-specific training data during the community prediction process, which effectively reduces interference from unrelated community structures. In comparison, CommunityAF employs a uniform training set for all nodes and lacks the capacity for personalized modeling.
@@ IV-B2 Comparison with semi-supervised community detection algorithms
TABLE V: Results of Bespoke, SEAL, CLARE, ProCom, and PPCS.
Datasets Metrics Bespoke-PPCS SEAL-PPCS CLARE-PPCS ProCom-PPCS
Bespoke PPCS SEAL PPCS CLARE PPCS ProCom PPCS
@@
As mentioned in Section I, the communities identified by Bespoke, SEAL, CLARE, and ProCom may not contain some given nodes. Therefore, we can only compare PPCS with these methods alone, using only the communities found by these methods that contain the given nodes. Table V shows the experimental results of Bespoke, SEAL, CLARE, ProCom, and PPCS on five datasets.
The results in Table V demonstrate that PPCS outperforms Bespoke, SEAL, CLARE, and ProCom on most datasets. In particular, PPCS achieves superior performance over SEAL on Amazon, DBLP, Twitter, and YouTube datasets, and its F-score and Jaccard values consistently surpass those of Bespoke, CLARE, and ProCom across all datasets. The relatively poor performance of Bespoke, SEAL, CLARE, and ProCom stems from their global search paradigms, which learn general community structures but fail to capture the structural characteristics of communities specific to the query node. As a result, the identified communities may not accurately reflect the ground-truth community structure of the query node. Additionally, SEAL achieves good performance on the LiveJournal dataset, primarily due to its generative adversarial network-based discriminator, which maintains high-quality community partitioning in the presence of large-scale and structurally diverse communities.
IV-C Time of execution.
@@
We compare the average runtime of PPCS with that of several semi-supervised community detection and semi-supervised community search methods on Amazon, DBLP, and Twitter, as shown in Figure 2 (a), (b), and (c). Each experiment uses 100 query nodes, with the average time recorded as the algorithm’s runtime.
The results show that PPCS is the fastest among all semi-supervised community search methods (SLSS, CommunityAF, and SLRL), achieving nearly 37% higher efficiency. The reasons are as follows: (1) The incremental GNN updates in CommunityAF are time-consuming. (2) PPCS avoids the costly matching of similar communities during community expansion, as required by SLSS, thus significantly reducing computational overhead. (3) Compared to SLRL, which performs more frequent community expansion during training, PPCS requires fewer expansion steps, thereby reducing the overall runtime.
In addition, PPCS is more runtime-efficient than semi-supervised community detection methods such as SEAL, CLARE, and ProCom. Specifically, PPCS differs from SEAL and CLARE in that its encoding process focuses on the local structural features of nodes. Moreover, during sample generation and prompt-driven fine-tuning, PPCS accesses only the neighbors of the query node, thereby reducing computational overhead. Although the runtime of ProCom is comparable to that of PPCS, the latter conducts community search based on the structure of the query node’s community, which incurs a higher computational cost but results in better community quality.
IV-D Ablation study.
@@
      Datasets       Metrics       w/o. NE       w/o. SG       w/o. PF       PPCS
@@
To evaluate the effectiveness of each component in PPCS, we conduct ablation studies on five datasets and design three simplified variants of the model: (1) “w/o. NE”: The node encoding component is removed, and node representations are directly derived from their original feature vectors; (2) “w/o. SG”: The sample generation component is removed, and the 2-ego subgraph of the given node is used as its initial community instead; (3) “w/o. PF”: The prompt-driven fine-tuning component is removed, and the initial community is directly treated as the final prediction community. We select 100 nodes as the given nodes to conduct experiments.
@@
Table VI presents the performance of these three simplified variants and the complete PPCS model across the datasets. Overall, PPCS consistently outperforms w/o. NE, w/o. SG, and w/o. PF, confirming the importance of the node encoding, sample generation, and prompt-driven fine-tuning components. More specifically, PPCS surpasses w/o. NE, indicating that the node encoding phase helps the model capture the features of local structures, which benefits downstream tasks. Compared to w/o. SG, the superior performance of PPCS demonstrates that high-quality initial communities lead to better training community sets, thereby enhancing the quality of prompt-based learning. Finally, compared to w/o. PF, PPCS achieves better performance in the metrics of F-score and Jaccard, demonstrating that using a training community set tailored to the given node as prompts can effectively enhance the quality of community prediction. Moreover, the prompt-based learning mechanism enables the model to capture the structural characteristics of the given node’s community.
IV-E Parameter study.
(b)          (a)
(c)           (b)
Figure 3: Results of PPCS parameter experiments.
@@
We analyze the effect of the hyperparameters $\gamma$ and $m$ on the performance of PPCS. $\gamma$ is the discount factor for computing cumulative rewards during sample generation, chosen from {0.25, 0.5, 0.75, 1}. $m$ is the number of samples used for prompt-driven fine-tuning, with values in {20, 30, 40, 50}. Specifically, we vary one parameter while keeping the others fixed and evaluate the changes in F-score across five datasets. The results are shown in Figure 3 (a) and (b).
@@
(1) Effect of $\gamma$: On Amazon, the performance of PPCS steadily improves as $\gamma$ increases from 0.25 to 0.75, then slightly declines when $\gamma$ reaches 1. For DBLP, Twitter, and YouTube, PPCS’s performance fluctuates slightly when $\gamma$ increases from 0.25 to 0.75, but peaks at $\gamma=\text{1}$. For LiveJournal, performance decreases as $\gamma$ increases from 0.25 to 0.75, and then slightly improves at $\gamma=\text{1}$. We set $\gamma=\text{1}$ for other experiments.
@@
(2) Effect of $m$: Even when only 20 similar samples are provided for each query node, PPCS achieves strong performance across all datasets. While performance generally improves with increasing $m$, the overall change from $m=\text{20}$ to $m=\text{50}$ remains modest. The default value of $m$ during the experiments is 20.
@@ V Conclusion
This paper proposes Pre-trained Prompt-driven Community Search (PPCS), the first attempt to apply the “pre-train, prompt” paradigm to semi-supervised community search. Unlike existing methods designed for semi-supervised community detection, PPCS focuses on identifying the community of a given node with improved accuracy and efficiency. The model integrates three core components: a node encoding component that learns node and community representations through local structural patterns using graph neural networks; a sample generation component that constructs training data by identifying an initial community for a given node and retrieving structurally similar communities; and a prompt-driven fine-tuning component that uses training samples as prompts to guide the final community prediction. Extensive experiments on five real-world datasets show that PPCS consistently outperforms baseline methods. Ablation studies validate the contribution of each component. In future work, we aim to further integrate the “pre-train, prompt” paradigm with large language models to enhance its applicability to more complex network analysis tasks.

### v2

@@ IV-A Settings
(1) An Unsupervised local community detection algorithm includes the M method.
@@
TABLE IV: Results of M, SLSS, ComAF, SLRL, and PPSL. The best results are highlighted in bold, and the second-best results are underlined.
Datasets Metrics M SLSS CommunityAF SLRL PPSL
@@
(2) Semi-supervised local community detection algorithms include SLSS, CommunityAF, and SLRL.
  • SLSS [24]: This method is the first semi-supervised local community detection algorithm, which proposes a graph kernel-based community structure similarity measurement.
  • CommunityAF [28]: This method is a semi-supervised local community detection model consisting of three components: a GNN-based feature extraction module, an autoregressive flow-based community generation module, and a scoring module.
@@ IV-B Results
IV-B1 Comparison with semi-supervised local community detection algorithms
We compared the results of M, SLSS, CommunityAF, SLRL, and PPSL. Table IV presents the experimental results for five datasets.
Experimental results indicate that PPSL outperforms M, SLSS, CommunityAF, and SLRL in terms of overall performance. On the Amazon, DBLP, Twitter, and YouTube datasets, PPSL achieves higher F-score and Jaccard values compared to the other methods, except the LiveJournal dataset, where its performance is slightly inferior. In general, semi-supervised local community detection algorithms (SLSS, CommunityAF, SLRL, and PPSL) tend to surpass the unsupervised method M, as they leverage known community information as prior knowledge, thereby guiding the model to identify community structures more accurately. However, the M method exhibits slightly better performance on the LiveJournal dataset because the community structure in this dataset aligns more closely with the modeling assumptions underlying the M method.
On most datasets other than LiveJournal, PPSL outperforms SLRL. This is primarily due to PPSL’s ability to learn structural characteristics of local structures through the node encoding component, which enhances the model’s generalization capacity. Additionally, PPSL incorporates a prompt mechanism to generate query-specific community predictions, thereby improving prediction accuracy. In contrast, SLRL performs relatively poorly on the LiveJournal dataset due to its inability to learn effective features with limited labeled data, whereas PPSL maintains strong predictive performance even with few samples. PPSL outperforms CommunityAF because it utilizes query-specific training data during the community prediction process, which effectively reduces interference from unrelated community structures. In comparison, CommunityAF employs a uniform training set for all nodes and lacks the capacity for personalized modeling.
@@ IV-B2 Comparison with semi-supervised community detection algorithms
TABLE V: Results of Bespoke, SEAL, CLARE, ProCom, and PPSL.
Datasets Metrics Bespoke-PPSL SEAL-PPSL CLARE-PPSL ProCom-PPSL
Bespoke PPSL SEAL PPSL CLARE PPSL ProCom PPSL
@@
As mentioned in Section I, the communities identified by Bespoke, SEAL, CLARE, and ProCom may not contain some given nodes. Therefore, we can only compare PPSL with these methods alone, using only the communities found by these methods that contain the given nodes. Table V shows the experimental results of Bespoke, SEAL, CLARE, ProCom, and PPSL on five datasets.
The results in Table V demonstrate that PPSL outperforms Bespoke, SEAL, CLARE, and ProCom on most datasets. In particular, PPSL achieves superior performance over SEAL on Amazon, DBLP, Twitter, and YouTube datasets, and its F-score and Jaccard values consistently surpass those of Bespoke, CLARE, and ProCom across all datasets. The relatively poor performance of Bespoke, SEAL, CLARE, and ProCom stems from their global search paradigms, which learn general community structures but fail to capture the structural characteristics of communities specific to the query node. As a result, the identified communities may not accurately reflect the ground-truth community structure of the query node. Additionally, SEAL achieves good performance on the LiveJournal dataset, primarily due to its generative adversarial network-based discriminator, which maintains high-quality community partitioning in the presence of large-scale and structurally diverse communities.
IV-C Time of execution
@@
We compare the average runtime of PPSL with that of several semi-supervised community detection and semi-supervised local community detection methods on Amazon, DBLP, and Twitter, as shown in Figure 2 (a), (b), and (c). Each experiment uses 100 query nodes, with the average time recorded as the algorithm’s runtime.
The results show that PPSL is the fastest among all semi-supervised local community detection methods (SLSS, CommunityAF, and SLRL), achieving nearly 37% higher efficiency. The reasons are as follows: (1) The incremental GNN updates in CommunityAF are time-consuming. (2) PPSL avoids the costly matching of similar communities during community expansion, as required by SLSS, thus significantly reducing computational overhead. (3) Compared to SLRL, which performs more frequent community expansion during training, PPSL requires fewer expansion steps, thereby reducing the overall runtime.
In addition, PPSL is more runtime-efficient than semi-supervised community detection methods such as SEAL, CLARE, and ProCom. Specifically, PPSL differs from SEAL and CLARE in that its encoding process focuses on the local structural features of nodes. Moreover, during sample generation and prompt-driven fine-tuning, PPSL accesses only the neighbors of the query node, thereby reducing computational overhead. Although the runtime of ProCom is comparable to that of PPSL, the latter conducts local community detection based on the structure of the query node’s community, which incurs a higher computational cost but results in better community quality.
IV-D Ablation study
@@
      Datasets       Metrics       w/o. NE       w/o. SG       w/o. PF       PPSL
@@
To evaluate the effectiveness of each component in PPSL, we conduct ablation studies on five datasets and design three simplified variants of the model: (1) “w/o. NE”: The node encoding component is removed, and node representations are directly derived from their original feature vectors; (2) “w/o. SG”: The sample generation component is removed, and the 2-ego subgraph of the given node is used as its initial community instead; (3) “w/o. PF”: The prompt-driven fine-tuning component is removed, and the initial community is directly treated as the final prediction community. We select 100 nodes as the given nodes to conduct experiments.
@@
Table VI presents the performance of these three simplified variants and the complete PPSL model across the datasets. Overall, PPSL consistently outperforms w/o. NE, w/o. SG, and w/o. PF, confirming the importance of the node encoding, sample generation, and prompt-driven fine-tuning components. More specifically, PPSL surpasses w/o. NE, indicating that the node encoding phase helps the model capture the features of local structures, which benefits downstream tasks. Compared to w/o. SG, the superior performance of PPSL demonstrates that high-quality initial communities lead to better training community sets, thereby enhancing the quality of prompt-based learning. Finally, compared to w/o. PF, PPSL achieves better performance in the metrics of F-score and Jaccard, demonstrating that using a training community set tailored to the given node as prompts can effectively enhance the quality of community prediction. Moreover, the prompt-based learning mechanism enables the model to capture the structural characteristics of the given node’s community.
IV-E Parameter study
(b)          (a) Effect of the discount factor $\gamma$
(c)           (b) Effect of sample numbers $m$
Figure 3: Results of PPSL parameter experiments.
@@ IV-E Parameter study
We analyze the effect of the hyperparameters $\gamma$ and $m$ on the performance of PPSL. $\gamma$ is the discount factor for computing cumulative rewards during sample generation, chosen from {0.25, 0.5, 0.75, 1}. $m$ is the number of samples used for prompt-driven fine-tuning, with values in {20, 30, 40, 50}. Specifically, we vary one parameter while keeping the others fixed and evaluate the changes in F-score across five datasets. The results are shown in Figure 3 (a) and (b).
@@
(1) Effect of $\gamma$: On Amazon, the performance of PPSL steadily improves as $\gamma$ increases from 0.25 to 0.75, then slightly declines when $\gamma$ reaches 1. For DBLP, Twitter, and YouTube, PPSL’s performance fluctuates slightly when $\gamma$ increases from 0.25 to 0.75, but peaks at $\gamma=\text{1}$. For LiveJournal, performance decreases as $\gamma$ increases from 0.25 to 0.75, and then slightly improves at $\gamma=\text{1}$. We set $\gamma=\text{1}$ for other experiments.
@@
(2) Effect of $m$: Even when only 20 similar samples are provided for each query node, PPSL achieves strong performance across all datasets. While performance generally improves with increasing $m$, the overall change from $m=\text{20}$ to $m=\text{50}$ remains modest. The default value of $m$ during the experiments is 20.
@@ V Conclusion
This paper proposes Pre-trained Prompt-driven Semi-supervised Local community detection (PPSL), the first attempt to apply the “pre-train, prompt” paradigm to semi-supervised local community detection. Unlike existing methods designed for semi-supervised community detection, PPSL focuses on identifying the community of a given node with improved accuracy and efficiency. The model integrates three core components: a node encoding component that learns node and community representations through local structural patterns using graph neural networks; a sample generation component that constructs training data by identifying an initial community for a given node and retrieving structurally similar communities; and a prompt-driven fine-tuning component that uses training samples as prompts to guide the final community prediction. Extensive experiments on five real-world datasets show that PPSL consistently outperforms baseline methods. Ablation studies validate the contribution of each component. In future work, we aim to further integrate the “pre-train, prompt” paradigm with large language models to enhance its applicability to more complex network analysis tasks.